i need to factor out completely 7x^2-30x+8 and say if its factorable or not factorable...could you solve it step by step .. please
is it 7x^2-28x-2x+8=0

Answer by ewatrrr(24785)

 
Hi,
7x^2-30x+8
factoring using the Grouping Method
7x^2-28x-2x+8=0
7x(x-4) -2(x-4) = 0
(7x - 2)(x-4) = 0
(7x - 2)= 0 x = 2/7
0r (x-4) = 0 x = 4
Checking Factoring Using FOIL
F First terms
O Outside terms -28x
I Inside terms -2x
L Last terms